OpenStudy (anonymous):

Medal for help Please ? The pair of points (6, y) and (10,-1) lie on a line with a slope 1/4. What is the value of y ?

OpenStudy (anonymous):

a.-5 b.-2 c.2 d.5

OpenStudy (anonymous):

c ****

OpenStudy (anonymous):

@thatfoolkasey

OpenStudy (anonymous):

-1 - Y1 _________ = 1/4 -1 - ( -2 ) = 1 the answer is B 10 - 6

OpenStudy (anonymous):

Gosh that - always gets me TY

OpenStudy (anonymous):

haha no problem you just have to practice thats why i told u what u have to do to solve and showed u how. algebra gets easier haha

OpenStudy (anonymous):